# SimpleMem: Efficient Lifelong Memory for LLM Agents

### v1

@@ 3.1 Experimental Setup
Benchmark Dataset. We utilize the LoCoMo benchmark (Maharana et al., 2024), which is specifically designed to test the limits of LLMs in processing long-term conversational dependencies. The dataset comprises conversation samples ranging from 200 to 400 turns, containing complex temporal shifts and interleaved topics. The evaluation set consists of 1,986 questions categorized into four distinct reasoning types: (1) Multi-Hop Reasoning: Questions requiring the synthesis of information from multiple disjoint turns (e.g., ‘‘Based on what X said last week and Y said today...’’); (2) Temporal Reasoning: Questions testing the model’s ability to understand event sequencing and absolute timelines (e.g., ‘‘Did X happen before Y?’’); (3) Open Domain: General knowledge questions grounded in the conversation context; (4) Single Hop: Direct retrieval tasks requiring exact matching of specific facts.
Baselines. We compare SimpleMem with representative memory-augmented systems: LoCoMo (Maharana et al., 2024), ReadAgent (Lee et al., 2024), MemoryBank (Zhong et al., 2024), MemGPT (Packer et al., 2023), A-Mem (Xu et al., 2025), LightMem (Fang et al., 2025), and Mem0 (Dev & Taranjeet, 2024).
@@
Implementation Details. For semantic structured compression, we use a sliding window of size $W=10$ and set the entropy-based significance threshold to $\tau=0.35$ to filter low-information interaction content. Memory indexing is implemented using LanceDB with a multi-view design: text-embedding-3-small (1536 dimensions) for dense semantic embeddings, BM25 for sparse lexical indexing, and SQL-based metadata storage for symbolic attributes. Recursive consolidation is triggered when the average pairwise semantic similarity within a memory cluster exceeds $\tau_{\text{cluster}}=0.85$. During retrieval, we employ adaptive query-aware retrieval, where the retrieval depth is dynamically adjusted based on estimated query complexity, ranging from $k_{\min}=3$ for simple lookups to $k_{\max}=20$ for complex reasoning queries.
@@
Evaluation Metrics. We report: F1 and BLEU-1 (accuracy), Adversarial Success Rate (robustness to distractors), and Token Cost (retrieval/latency efficiency). LongMemEval-S uses its standard accuracy-style metric.
3.2 Main Results and Analysis
We evaluate SimpleMem across a diverse set of LLMs, ranging from high-capability proprietary models (GPT-4o series) to efficient open-source models (Qwen series). Tables 1 and 2 present the detailed performance comparison on the LoCoMo benchmark.
Performance on High-Capability Models. As shown in Table 1, SimpleMem consistently outperforms existing memory systems across all evaluated models. On GPT-4.1-mini, SimpleMem achieves an Average F1 of 43.24, establishing a significant margin over the strongest baseline, Mem0 (34.20), and surpassing the full-context baseline (LoCoMo, 18.70) by over 24 points. Notable gains are observed in Temporal Reasoning, where SimpleMem scores 58.62 F1 compared to Mem0’s 48.91, demonstrating the effectiveness of our Semantic Structured Compression in resolving complex timelines. Similarly, on the flagship GPT-4o, SimpleMem maintains its lead with an Average F1 of 39.06, outperforming Mem0 (36.09) and A-Mem (33.45). These results confirm that Recursive Consolidation mechanism effectively distills high-density knowledge, enabling even smaller models equipped with SimpleMem to outperform larger models using traditional memory systems.
@@ 3.2 Main Results and Analysis
Token Efficiency. A key strength of SimpleMem lies in its inference-time efficiency. As reported in the rightmost columns of Tables 1 and 2, full-context approaches such as LoCoMo and MemGPT consume approximately 16,900 tokens per query. In contrast, SimpleMem reduces token usage by roughly $30\times$, averaging 530–580 tokens per query. Furthermore, compared to optimized retrieval baselines like Mem0 ($\sim$980 tokens) and A-Mem ($\sim$1,200+ tokens), SimpleMem reduces token usage by 40-50% while delivering superior accuracy. For instance, on GPT-4.1-mini, SimpleMem uses only 531 tokens to achieve state-of-the-art performance, whereas ReadAgent consumes more (643 tokens) but achieves far lower accuracy (7.16 F1). This validates the efficacy of our Entropy-based Filtering and Adaptive Pruning, which strictly control context bandwidth without sacrificing information density.
@@
Robustness Across Task Types. Breaking down performance by task, SimpleMem demonstrates balanced capabilities. In SingleHop QA, it consistently leads (e.g., 51.12 F1 on GPT-4.1-mini), proving precision in factual retrieval. In complex MultiHop scenarios, SimpleMem significantly outperforms Mem0 and LightMem on GPT-4.1-mini, indicating that our Molecular Representations successfully bridge disconnected facts, enabling deep reasoning without the need for expensive iterative retrieval loops.
@@ 3.4 Ablation Study
To verify the claims that specific cognitive mechanisms correspond to computational gains, we conducted a component-wise ablation study using the GPT-4.1-mini backend. We investigate the contribution of three key components: (1) Semantic Structured Compression , (2) Recursive Consolidation, and (3) Adaptive Query-Aware Retrieval. The results are summarized in Table 4.
@@
   w/o Atomization 34.20 ($\downarrow$21.3%) 25.40 ($\downarrow$56.7%) 17.50 ($\downarrow$11.4%) 48.05 ($\downarrow$6.0%) 31.29 ($\downarrow$27.6%)
   w/o Consolidation 29.85 ($\downarrow$31.3%) 55.10 ($\downarrow$6.0%) 18.20 ($\downarrow$7.9%) 49.80 ($\downarrow$2.6%) 38.24 ($\downarrow$11.6%)
   w/o Adaptive Pruning 38.60 ($\downarrow$11.2%) 56.80 ($\downarrow$3.1%) 14.50 ($\downarrow$26.6%) 41.20 ($\downarrow$19.4%) 37.78 ($\downarrow$12.6%)
@@
Impact of Recursive Consolidation. Disabling the background consolidation process results in a 31.3% decrease in multi-hop reasoning performance. Without consolidating related memory units into higher-level abstract representations, the system must retrieve a larger number of fragmented entries during reasoning. This fragmentation increases context redundancy and exhausts the available context window in complex queries, demonstrating that recursive consolidation is essential for synthesizing dispersed evidence into compact and informative representations.
Impact of Adaptive Query-Aware Retrieval. Removing the adaptive retrieval mechanism and reverting to fixed-depth retrieval primarily degrades performance on open-domain and single-hop tasks, with drops of 26.6% and 19.4%, respectively. In the absence of query-aware adjustment, the system either retrieves insufficient context for entity-specific queries or introduces excessive irrelevant information for simple queries. These results highlight the importance of dynamically modulating retrieval scope to balance relevance and efficiency during inference.
@@ 3.5 Case Study: Long-Term Temporal Grounding
Precise Retrieval. When queried about Sarah’s past artworks, the adaptive retrieval mechanism combines semantic relevance with symbolic constraints to exclude unrelated activities and retrieve only temporally valid entries. The system correctly identifies relevant paintings while ignoring semantically related but irrelevant topics. This example demonstrates how structured compression, temporal normalization, and adaptive retrieval jointly enable reliable long-term reasoning under extended interaction histories.
@@ 4 Related Work
Memory Systems for LLM Agents. Recent approaches manage memory through virtual context or structured representations. Virtual context methods, including MemGPT (Packer et al., 2023), MemoryOS (Kang et al., 2025), and SCM (Wang et al., 2023), extend interaction length via paging or stream-based controllers (Wang et al., 2024b) but typically store raw conversation logs, leading to redundancy and increasing processing costs. In parallel, structured and graph-based systems, such as MemoryBank (Zhong et al., 2024), Mem0 (Dev & Taranjeet, 2024), Zep (Rasmussen et al., 2025), A-Mem (Xu et al., 2025), and O-Mem (Wang et al., 2025), impose structural priors to improve coherence but still rely on raw or minimally processed text, preserving referential and temporal ambiguities that degrade long-term retrieval. In contrast, SimpleMem adopts a semantic compression mechanism that converts dialogue into independent, self-contained facts, explicitly resolving referential and temporal ambiguities prior to storage.
Context Management and Retrieval Efficiency. Beyond memory storage, efficient access to historical information remains a core challenge. Existing approaches primarily rely on either long-context models or retrieval-augmented generation (RAG). Although recent LLMs support extended context windows (OpenAI, 2025; Deepmind, 2025; Anthropic, 2025), and prompt compression methods aim to reduce costs (Jiang et al., 2023a; Liskavetsky et al., 2025), empirical studies reveal the “Lost-in-the-Middle” effect (Liu et al., 2023; Kuratov et al., 2024), where reasoning performance degrades as context length increases, alongside prohibitive computational overhead for lifelong agents. RAG-based methods (Lewis et al., 2020; Asai et al., 2023; Jiang et al., 2023b), including structurally enhanced variants such as GraphRAG (Edge et al., 2024; Zhao et al., 2025) and LightRAG (Guo et al., 2024), decouple memory from inference but are largely optimized for static knowledge bases, limiting their effectiveness for dynamic, time-sensitive episodic memory. In contrast, SimpleMem improves retrieval efficiency through Adaptive Pruning and Retrieval, jointly leveraging semantic, lexical, and metadata signals to enable precise filtering by entities and timestamps, while dynamically adjusting retrieval depth based on query complexity to minimize token usage.
@@ 5 Conclusion
We introduce SimpleMem, an efficient memory architecture governed by the principle of Semantic Lossless Compression. By reimagining memory as a metabolic process, SimpleMem implements a dynamic continuum: Semantic Structured Compression to filter noise at the source, Recursive Consolidation to evolve fragmented facts into high-order molecular insights, and Adaptive Spatial Pruning to dynamically modulate retrieval bandwidth. Empirical evaluation on the LoCoMo benchmark demonstrates the effectiveness and efficiency of SimpleMem.
@@ Acknowledgement
This work is partially supported by Amazon Research Award, Cisco Faculty Research Award, and Coefficient Giving.
@@ Appendix B Extended Implementation Details and Experiments
B.1 Hyperparameter Configuration
@@
B.2 Hyperparameter Sensitivity Analysis
@@
Stage 1: Semantic Structured Compression Window Size ($W$) 10 turns
Sliding Stride 5 turns (50% overlap)
Information Threshold ($\tau$) 0.35 (filters low-information interaction content)
Model Backend gpt-4o-mini (temperature = 0.0)
Coreference Scope Current window with up to two preceding turns
@@
Stage 2: Recursive Consolidation Embedding Model text-embedding-3-small (1536 dimensions)
Consolidation Threshold ($\tau_{\text{cluster}}$) 0.85 (triggers abstraction over related memory units)
Temporal Decay ($\lambda$) 0.1 (controls temporal influence during consolidation)
@@
Stage 3: Adaptive Retrieval Query Complexity Estimator gpt-4o-mini (classification head)
Retrieval Range $k\in[3,20]$
Minimum Depth ($k_{\min}$) 3 (symbolic and abstract-level retrieval)
Maximum Depth ($k_{\max}$) 20 (expanded semantic retrieval)

### v2

@@ 3.1 Experimental Setup
Benchmark Dataset. We evaluate performance on the LoCoMo (Maharana et al., 2024) and LongMemEval-S (Wu et al., 2024) benchmarks. Brief descriptions are provided below, with additional details in Appendix B.
LoCoMo (Maharana et al., 2024) is specifically designed to test the limits of LLMs in processing long-term conversational dependencies. The dataset comprises conversation samples ranging from 200 to 400 turns, containing complex temporal shifts and interleaved topics. The evaluation set consists of 1,986 questions categorized into four distinct reasoning types: (1) multi-hop reasoning; (2) temporal reasoning; (3) open Domain; (4) single hop.
LongMemEval-S (Wu et al., 2024) features extreme context lengths that pose severe challenges for memory systems. Unlike standard benchmarks, it requires precise answer localization across multiple sub-categories (e.g., temporal events, user preferences) within exceptionally long interaction histories. We use gpt-4.1-mini to evaluate answer correctness against ground-truth references, labeling responses as Correct or Wrong based on semantic and temporal alignment. The full evaluation prompt is provided in Appendix A.4.
Baselines. We compare SimpleMem with representative memory-augmented systems: LoCoMo (Maharana et al., 2024), ReadAgent (Lee et al., 2024), MemoryBank (Zhong et al., 2024), MemGPT (Packer et al., 2023), A-Mem (Xu et al., 2025), LightMem (Fang et al., 2025), and Mem0 (Dev and Taranjeet, 2024).
@@
Implementation Details. For semantic structured compression, we use a sliding window of size $W=10$. Memory indexing is implemented using LanceDB with a multi-view design: Qwen3-embedding-0.6b (1024 dimensions) for dense semantic embeddings, BM25 for sparse lexical indexing, and SQL-based metadata storage for symbolic attributes. During retrieval, we employ adaptive query-aware retrieval, where the retrieval depth is dynamically adjusted based on estimated query complexity, ranging from $k_{\min}=3$ for simple lookups to $k_{\max}=20$ for complex reasoning queries.
@@
Evaluation Metrics. For LoCoMo, we report F1 and BLEU-1 (accuracy), Adversarial Success Rate (robustness to distractors), and Token Cost (retrieval efficiency). For LongMemEval-S, we use its standard accuracy-style metric.
3.2 Results and Analysis
Tables 1 and 3 present detailed performance comparisons on the LoCoMo benchmark across different model scales, while Table 2 reports results on LongMemEval-S.
Performance on High-Capability Models. Across LoCoMo and LongMemEval-S, SimpleMem consistently outperforms existing memory systems across model scales, achieving strong and robust gains in accuracy.
On LoCoMo (Table 1), SimpleMem leads all baselines. Using GPT-4.1-mini, it achieves an Average F1 of 43.24, substantially exceeding Mem0 (34.20) and the full-context baseline (18.70). The largest gains are observed in Temporal Reasoning, where SimpleMem reaches 58.62 F1 compared to 48.91 for Mem0, underscoring the effectiveness of semantic structured compression in resolving complex temporal dependencies. These improvements persist at larger scales: on GPT-4o, SimpleMem attains the highest Average F1 (39.06), outperforming Mem0 (36.09) and A-Mem (33.45).
A task-level breakdown on LoCoMo further highlights the balanced capabilities of SimpleMem. In SingleHop QA, SimpleMem consistently achieves the best performance (e.g., 51.12 F1 on GPT-4.1-mini), demonstrating precise factual retrieval. In more challenging MultiHop settings, SimpleMem significantly outperforms Mem0 and LightMem on GPT-4.1-mini, indicating its ability to bridge disconnected facts and support deep reasoning without relying on expensive iterative retrieval loops.
Results on LongMemEval-S (Table 2) further demonstrate the robustness of SimpleMem under extreme context lengths. Using the gpt-4.1-mini backbone, SimpleMem achieves the highest average accuracy of 72.40%, outperforming LightMem (69.20%) and substantially exceeding Mem0 (56.40%) and the full-context baseline (35.40%). Gains are most pronounced in the challenging Multi-Session category, where SimpleMem attains 60.92% accuracy, compared to 47.37% for LightMem and 30.08% for full-context, highlighting its effectiveness in cross-session information integration under severe context constraints.
When scaled to the more capable gpt-4.1 backbone, SimpleMem maintains its state-of-the-art performance with an average accuracy of 77.60%. It is particularly strong in the Single-Session User task (98.57%), demonstrating near-perfect recall of immediate user-provided details. While LightMem exhibits strong performance on temporally specific queries, SimpleMem offers a more balanced profile: it avoids catastrophic failures in assistant-focused recall (where LightMem drops to 21.43% on the mini model) while maintaining high accuracy in complex multi-session retrieval. This balance suggests that SimpleMem’s structured indexing strategy effectively disentangles episodic noise from salient facts, providing a reliable memory substrate for long-term interaction.
@@ 3.2 Results and Analysis
Token Efficiency. A key strength of SimpleMem lies in its inference-time efficiency. As reported in the rightmost columns of Tables 1 and 3, full-context approaches such as LoCoMo and MemGPT consume approximately 16,900 tokens per query. In contrast, SimpleMem reduces token usage by roughly $30\times$, averaging 530-580 tokens per query. Furthermore, compared to optimized retrieval baselines like Mem0 ($\sim$980 tokens) and A-Mem ($\sim$1,200+ tokens), SimpleMem reduces token usage by 40-50% while delivering superior accuracy. For instance, on GPT-4.1-mini, SimpleMem uses only 531 tokens to achieve state-of-the-art performance, whereas ReadAgent consumes more (643 tokens) but achieves far lower accuracy (7.16 F1). This validates the effectiveness of SimpleMem in strictly controlling context bandwidth without sacrificing information density.
@@ 3.4 Ablation Study
In addition, we conduct an ablation study using the GPT-4.1-mini backend. We investigate the contribution of three key components. The results are summarized in Table 5.
@@
   w/o Semantic Compression 34.20 ($\downarrow$21.3%) 25.40 ($\downarrow$56.7%) 17.50 ($\downarrow$11.4%) 48.05 ($\downarrow$6.0%) 31.29 ($\downarrow$27.6%)
   w/o Online Synthesis 29.85 ($\downarrow$31.3%) 55.10 ($\downarrow$6.0%) 18.20 ($\downarrow$7.9%) 49.80 ($\downarrow$2.6%) 38.24 ($\downarrow$11.6%)
   w/o Intent-Aware Retrieval 38.60 ($\downarrow$11.2%) 56.80 ($\downarrow$3.1%) 14.50 ($\downarrow$26.6%) 41.20 ($\downarrow$19.4%) 37.78 ($\downarrow$12.6%)
@@
Impact of Online Semantic Synthesis. Disabling online semantic synthesis results in a 31.3% decrease in multi-hop reasoning performance. Without on-the-fly consolidation during the write phase, semantically related facts accumulate as fragmented entries, forcing the retriever to assemble dispersed evidence at query time. This fragmentation inflates contextual redundancy and rapidly exhausts the available context window in complex queries. The observed degradation demonstrates that proactive, intra-session synthesis is essential for maintaining a compact and semantically coherent memory topology, and for transforming local observations into reusable, high-density abstractions.
Intent-Aware Retrieval Planning. Removing intent-aware retrieval planning and reverting to a fixed-depth retrieval strategy primarily degrades performance on open-domain and single-hop tasks, with drops of 26.6% and 19.4%, respectively. In the absence of query-aware adjustment, the system either retrieves insufficient context for entity-specific queries or introduces excessive irrelevant information for simple queries. These results highlight the importance of dynamically modulating retrieval scope to balance relevance and efficiency during inference.
@@ 3.5 Case Study: Long-Term Temporal Grounding
Precise Retrieval. When queried about Sarah’s past artworks, the intent-aware retrieval planner infers both the semantic focus (art-related activities) and the temporal constraints implied by the query. The system then performs parallel multi-view retrieval, combining semantic similarity with symbolic filtering to exclude unrelated activities and return only temporally valid entries. This example demonstrates how structured compression, temporal normalization, and adaptive retrieval jointly enable reliable long-term reasoning under extended interaction histories.
@@ 4 Related Work
Memory Systems for LLM Agents. Recent approaches manage memory through virtual context or structured representations. Virtual context methods, including MemGPT (Packer et al., 2023), MemoryOS (Kang et al., 2025), and SCM (Wang et al., 2023), extend interaction length via paging or stream-based controllers (Wang et al., 2024b) but typically store raw conversation logs, leading to redundancy and increasing processing costs. In parallel, structured and graph-based systems, such as MemoryBank (Zhong et al., 2024), Mem0 (Dev and Taranjeet, 2024), Zep (Rasmussen et al., 2025), A-Mem (Xu et al., 2025), and O-Mem (Wang et al., 2025), impose structural priors to improve coherence but still rely on raw or minimally processed text, preserving referential and temporal ambiguities that degrade long-term retrieval. In contrast, SimpleMem adopts a semantic compression mechanism that converts dialogue into independent, self-contained facts, explicitly resolving referential and temporal ambiguities prior to storage.
Context Management and Retrieval Efficiency. Beyond memory storage, efficient access to historical information remains a core challenge. Existing approaches primarily rely on either long-context models or retrieval-augmented generation (RAG). Although recent LLMs support extended context windows (OpenAI, 2025; Deepmind, 2025; Anthropic, 2025), and prompt compression methods aim to reduce costs (Jiang et al., 2023a; Liskavetsky and others, 2025), empirical studies reveal the “Lost-in-the-Middle” effect (Liu et al., 2023; Kuratov and others, 2024), where reasoning performance degrades as context length increases, alongside prohibitive computational overhead for lifelong agents. RAG-based methods (Lewis et al., 2020; Asai et al., 2023; Jiang et al., 2023b), including structurally enhanced variants such as GraphRAG (Edge et al., 2024; Zhao et al., 2025) and LightRAG (Guo et al., 2024), decouple memory from inference but are largely optimized for static knowledge bases, limiting their effectiveness for dynamic, time-sensitive episodic memory. In contrast, SimpleMem improves retrieval efficiency through Intent-Aware Retrieval Planning, jointly leveraging semantic, lexical, and symbolic signals to construct query-specific retrieval plans and dynamically adapt the retrieval budget, achieving token-efficient reasoning under constrained context budgets.
@@ 5 Conclusion
We introduce SimpleMem, an efficient agent memory architecture grounded in the principle of semantic lossless compression. By treating memory as an active process rather than passive storage, SimpleMem integrates Semantic Structured Compression to filter noise at the source, Online Semantic Synthesis to consolidate fragmented observations during writing, and Intent-Aware Retrieval Planning to dynamically adapt retrieval scope. Empirical evaluation on the LoCoMo and LongMemEval-S benchmark demonstrates the effectiveness and efficiency of our method.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Use of AI Assistants
We acknowledge the use of AI assistants in the writing of this paper, mostly for formatting and polishing the text. All the generated content is reviewed by the authors and necessary corrections are made.
@@ Acknowledgement
This work is partially supported by Amazon Research Award, Cisco Faculty Research Award, and Coefficient Giving. This work is also partly supported by the National Center for Transportation Cybersecurity and Resiliency (TraCR) (a U.S. Department of Transportation National University Transportation Center) headquartered at Clemson University, Clemson, South Carolina, USA (USDOT Grant #69A3552344812). Any opinions, findings, conclusions, and recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of TraCR, and the U.S. Government assumes no liability for the contents or use thereof.
@@ 4. Faithfulness:
A.4 LongMemEval Evaluation Prompt
For the LongMemEval benchmark, we employed gpt-4.1-mini as the judge to evaluate the correctness of the agent’s responses. The prompt strictly instructs the judge to focus on semantic and temporal consistency rather than exact string matching. The specific prompt template used is provided below:
Listing 4: LLM-as-a-Judge Evaluation Prompt.
Your task is to label an answer to a question as CORRECT or WRONG’.
You will be given the following data:
(1) a question (posed by one user to another user),
(2) a gold (ground truth) answer,
(3) a generated answer
which you will score as CORRECT/WRONG.
The point of the question is to ask about something one user should know about the other user based on their prior conversations.
The gold answer will usually be a concise and short answer that includes the referenced topic, for example:
Question: Do you remember what I got the last time I went to Hawaii?
Gold answer: A shell necklace
The generated answer might be much longer, but you should be generous with your grading - as long as it touches on the same topic as the gold answer, it should be counted as CORRECT.
For time related questions, the gold answer will be a specific date, month, year, etc. The generated answer might be much longer or use relative time references (like "last Tuesday" or "next month"), but you should be generous with your grading - as long as it refers to the same date or time period as the gold answer, it should be counted as CORRECT. Even if the format differs (e.g., "May 7th" vs "7 May"), consider it CORRECT if its the same date.
Now its time for the real question:
Question: {question}
Gold answer: {gold_answer}
Generated answer: {generated_answer}
First, provide a short (one sentence) explanation of your reasoning, then finish with CORRECT or WRONG.
Do NOT include both CORRECT and WRONG in your response, or it will break the evaluation script.
Just return the label CORRECT or WRONG in a json format with the key as "label".
@@ Appendix B Extended Implementation Details and Experiments
B.1 Dataset Description
LoCoMo (Maharana et al., 2024) is specifically designed to test the limits of LLMs in processing long-term conversational dependencies. The dataset comprises conversation samples ranging from 200 to 400 turns, containing complex temporal shifts and interleaved topics. The evaluation set consists of 1,986 questions categorized into four distinct reasoning types: (1) Multi-Hop Reasoning: Questions requiring the synthesis of information from multiple disjoint turns (e.g., ‘‘Based on what X said last week and Y said today...’’); (2) Temporal Reasoning: Questions testing the model’s ability to understand event sequencing and absolute timelines (e.g., ‘‘Did X happen before Y?’’); (3) Open Domain: General knowledge questions grounded in the conversation context; (4) Single Hop: Direct retrieval tasks requiring exact matching of specific facts.
LongMemEval-S benchmark. The defining characteristic of this dataset is its extreme context length, which poses a unique and severe challenge for memory systems. Unlike standard benchmarks, LongMemEval-S requires the system to precisely locate specific answers across various sub-categories (e.g., temporal events, user preferences) within an exceptionally long interaction history. This massive search space significantly escalates the difficulty of retrieval and localization, serving as a rigorous stress test for the system’s precision. We utilized an LLM-as-a-judge protocol (using gpt-4.1-mini) to score the correctness of generated answers against ground-truth references, categorizing responses as either Correct or Wrong based on semantic and temporal alignment. The full evaluation prompt is provided in Appendix A.4.
B.2 Hyperparameter Configuration
@@
B.3 Hyperparameter Sensitivity Analysis
@@
Stage 1: Semantic Structured Compression Window Size ($W$) 20 turns
Sliding Stride 5 turns (25% overlap)
Model Backend gpt-4.1-mini (temperature = 0.0)
@@
Stage 2: Online Semantic Synthesis Embedding Model Qwen3-embedding-0.6b (1024 dimensions)
@@
Stage 3: Intent-Aware Retrieval Planning Query Complexity Estimator gpt-4.1-mini
Retrieval Range [3, 20]
Minimum Depth 1
Maximum Depth 20